# Synergizing Large Language Models and Task-specific Models for Time Series Anomaly Detection

### v3

@@ Abstract
Unsupervised Anomaly Detection, Time Series, Large Language Model, Task-specific Small Model, Model Collaboration
@@ 1 Introduction
Recently, methods based on Large Language Models (LLMs) have demonstrated the capablility of reading professional documents, effectively acquiring human knowledge to guide relevant tasks. However, they are often insensitive to the value fluctuations in time series data, and their natural language processing-based representations do not align well with the characteristics of time series data (Jin et al., 2024). In contrast, task-specific small scaled models, such as task-specific anomaly detection models (TSADM), typically lack the capabilities of reading professional documents and extracting expert knowledge. However, these models are specifically designed to learn patterns for anomaly detection and often exhibit superior performance when applied to well-matched anomaly detection datasets (Zhou et al., 2023). Despite their strengths, TSADMs also have notable limitations. First, for complex systems, such as cloud service system and flight systems, which require lots of expert knowledge to detect and diagnose anomalies accurately, the anomaly assumptions of unsupervised anomaly detection methods can be inconsistent with these scenarios (Zha et al., 2020). For instance, anomalies identified based on the voting mechanism of redundant aircraft channels (Rice & Mccorkle, 1979) may differ from the outlier assumptions utilized by unsupervised anomaly detection methods (Blázquez-García et al., 2021). Thus, researchers need to modify TSADMs to incorporate the background expertise to achieve optimal performance. For instance, anomaly detection methods tailored for cloud service monitoring (Ma et al., 2021; Chen et al., 2024b) or aircraft monitoring (e Silva & Murcca, 2023) have been modified to suit these specific contexts. Second, in many practical scenarios, such as aircraft monitoring (Nanduri & Sherry, 2016), it is not feasible to collect system monitoring data for all possible flight conditions. These conditions often represent distinct distributions and normal patterns in the monitoring data. This insufficient coverage of comprehensive data poses a significant challenge, greatly impeding the performance of TSADM.
@@
(a) The performance of CoLLaTe, LLM and TSADM
(b) Misalignment between anomaly scores of the LLM and TSADM
Figure 1: (a) The anomaly scores of LLM, TSADM, and CoLLaTe. (b) The normalized anomaly score distribution of TSADM and the fitted curve of anomaly scores from the LLM on Mustang (Amvrosiadis et al., 2018).
@@
2 Preliminaries
Given a TSADM and an LLM, they assign an anomaly score and a representation to each time slot of a time series $X$, as shown in Eq. 1–Eq. 2, where $\dot{s}$ and $S$ are anomaly scores given by the TSADM and the LLM, respectively, and $R$ is the representation of time slots. $X_{i}$ is the $i^{th}$ time slot of $X$. Similarly, $\dot{s}_{i}$ and $S_{i}$ are anomaly scores of the $i^{th}$ time slot, respectively. $y_{i}$ equals $1$ when $X_{i}$ is anomalous, and it equals 00 otherwise. CoLLaTe aims to utilize the representation $R$ as a condition to synthesize the anomaly scores from the two models, as shown in Eq. 3. By doing so, CoLLaTe produces a collated anomaly score $\hat{S}$.
$\dot{s},R=\operatorname{SmallModel}(X)$ (1)
$S=\operatorname{LLM}(X)$ (2)
$\hat{S}=\operatorname{CoLLaTe}(\dot{s},S;R)$ (3)
@@ 3.1 Overview
The architecture of CoLLaTe is depicted in Fig. 2. The anomaly scores, $\dot{s}$ and $S$, are generated by a TSADM and an LLM, respectively, where a set-up pitch prompt is utilized to embed expert knowledge from professional document to the LLM (Appendix. A.11). The TSADM is an anomaly attention (Xu et al., 2022) based anomaly detection model, whose architecture is illustrated in Section. A.9 in detail. These anomaly scores, $\dot{s}$ and $S$, typically represent different interpretations of the same score, as the LLM and the TSADM can assign the same score to denote varying degrees of anomaly. To address this, we introduce an alignment module to align the different interpretations of the anomaly scores between the LLM and the TSADM. Following alignment, we employ a conditional network to synthesize the aligned scores from the TSADM and anomaly scores from the LLM, using the data representation as a condition. This process yields the collated anomaly score, $\hat{S}$. Additionally, we propose a collaborative loss function that effectively leverages the complementary strengths of the LLM and the TSADM. Notably, we prove this loss function can mitigate error accumulation during the collaborative process between the LLM and the TSADM mathematically and experimentally.
@@ 3.2 Alignment Module
TSADM and large language models (LLMs) often interpret anomaly scores differently. In other words, they may assign the same anomaly score to represent varying degrees of anomaly severity. This discrepancy arises due to two main factors: value range misalignment and distribution misalignment. The value range misalignment is shown in Fig. 3(b), where we shift the fitted curve of LLM anomaly scores to the histogram of TSADM anomaly scores. Most reconstruction-based TSADMs use reconstruction error as the anomaly score, resulting in a highly variable range of scores. In contrast, the anomaly scores produced by LLMs are typically constrained to the range $[0,1]$. This problem can be easily solved by min-max normalization and we obtain the scaled anomaly scores of TSADM, $s$. This process is illustrated in detail in Appendix. A.12.
Even after scaling the anomaly scores of the TSADM, differences in the distribution of anomaly scores can still lead to inconsistent interpretations. As shown in Fig. 3(a), we use a half-Gaussian distribution, which will be elaborated in the following, to fit the histogram of LLM anomaly scores. Then, we shift the fitted curve to the histogram of the anomaly scores from the TSADM in Fig. 1(b), where we observe that most LLM anomaly scores fall below 0.3, whereas the TSADM’s anomaly scores are concentrated in the range of [0.3, 0.4]. As a result, an anomaly score of 0.4 might indicate moderate anomaly risks for the LLM, while the TSADM may interpret it as normal. In such cases, direct collaboration between the TSADM and the LLM without proper alignment could misrepresent their intended outputs. For example, if the LLM assigns a score of 0.4 to a time slot, the TSADM might incorrectly assume the LLM considers the time slot normal. To address this challenge, we propose an Alignment Module to align their interpretations and ensure effective collaboration.
@@
Subsequently, we use a mapping function $\mathcal{M}(s)$ to map $s\in\{s_{i}\}_{i=0}^{n}$ from the TSADM to a new set, whose distribution should approach $f(S)$. To achieve this, we divide the value range of anomaly scores, $[0,1]$, into N bins. We use $c_{i}$ to denote how many mapped scores $\mathcal{M}(s)$ fall into the $i^{th}$ bin. Then, we minimize cross entropy to decrease the distance between $f(S)$ and the distribution mapped set as Eq. 5 shows, where $\hat{\mu},\hat{\sigma}$ are the mean and standard variance of distribution $f(S)$ and $\mu_{M}=\frac{1}{n}\sum_{k=1}^{n}\mathcal{M}(s_{k})$. One problem with Eq. 5 is that the counting process of $c_{i}$ is non-differentiable and could not be used in backforward propagation.
@@
To solve the above problem, we propose Theorem 1 and we use Eq. 6 as our loss function instead.
Theorem 1. When $N$ approaches infinite, Eq. 5 can be transformed to Eq. 6, which is differentiable. $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are hyperparameters.
The proof of Theorem 1 is given in Appendix. A.1.
$\begin{split}\min_{\mathcal{M}}.\mathcal{L}_{a}=&-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}\\ &+\hat{\lambda}_{2}(\frac{1}{n-1}\sum_{i=1}^{n}(\mathcal{M}(s_{i})-\mu_{M})^{2}-\hat{\sigma}^{2})^{2}\end{split}$ (6)
(a) Anomaly score distribution of large language model
(b) Misalignment between original anomaly scores of LLM and TSADM
Figure 3: (a) shows the histogram of the anomaly score of LLM on the Mustang (Amvrosiadis et al., 2018) dataset and the fitted curve of the anomaly score. (b) The original anomaly score distribution of TSADM for anomaly detection and the fitted curve of anomaly scores from the LLM on Mustang.
@@ 3.3 Collaborative Loss
Many anomaly detection studies categorize anomalies into contextual anomalies and point anomalies (Schmidl et al., 2022; Tuli et al., 2022). We conducted experiments on two aircraft monitoring datasets, which are detailed in Appendix A.6. The results, presented in Tab. 1, demonstrate the complementary performance of the TSADM and the LLM in detecting contextual and point anomalies. Specifically, the F1 score for contextual anomalies is higher for the TSADM compared to GPT-4, while GPT-4 achieves a better F1 score for point anomalies than the TSADM. According to this insight, we propose a collaborative loss function that adaptively optimizes the collated anomaly score. Specifically, we begin by dividing the time series into patches (Nie et al., 2023). For each time slot, we compute the average distance between the present time slot and all other time slots within the same patch, denoted as $D_{intra}$. Next, we calculate the average distance between the patch of the current time slot and all other patches, denoted as $D_{inter}$. Since anomalous time slots differ significantly from normal ones, while normal time slots tend to share similarities, the metrics $D_{intra}$ and $D_{inter}$ behave differently for point and contextual anomalies. For point anomalies, $D_{intra}$ should be large because the point anomaly stands out distinctly from the surrounding normal time slots. Conversely, $D_{inter}$ for point anomalies should be smaller, as the normal time slots within the current patch resemble those in other patches, reducing the average difference. For contextual anomalies, which typically involve a sustained deviation, the anomalies within an anomalous event are similar to each other but differ significantly from normal time slots. As a result, $D_{intra}$ should be small, while $D_{inter}$ should be large. Considering LLM can handle point anomaly better and TSADM can handle contextual anomaly better, we design the collaborative loss function in Eq. 8, where $\mathcal{L}(a,b)$ represents a loss function that measures the distance between $a$ and $b$.
@@
Assumption 3. $\mathcal{L}_{\beta}^{*}$ in Eq. 9 is convex and $L$-Lipschitz continuous.
@@
From Theorem 2, we observe that when MSE is used as $\mathcal{L}$, the expected prediction error of the conditional network exceeds the weighted sum of the expected prediction errors of the TSADM and LLM. Given that the weights $\frac{D_{intra}}{D_{intra}+D_{inter}}$ and $\frac{D_{inter}}{D_{intra}+D_{inter}}$ are both greater than 0 and sum to 1, it indicates that the collaborative error, when using MSE as $L_{(a,b)}$, consistently exceeds the smaller of the errors produced by the LLM and TSADM. This phenomenon reflects error accumulation and is further validated experimentally, as discussed in detail in Section 4.4.
To solve this problem we propose the collaborative loss function as shown in Eq. 10, where $\hat{S}_{i}$, $s_{i}$ and $S_{i}$ are anomaly scores of $i^{t}h$ sample output by the conditional network, TSADM and LLM respectively. In Lemma 1, we prove that using the collaborative loss function is approximate to using the loss function in Eq. 9. This equivalence shows that Eq. 10 prevents error accumulation between the TSADM and LLM, as Eq.9 directly minimizes the difference between the collated anomaly scores of two samples and the ground truth. Additionally, we experimentally confirm that Eq. 10 effectively avoids error accumulation, as discussed in Section. 4.4. Moreover, in Theorem 3, we prove that the optimal solution of Eq. 9 satisfies two key properties. Property 1 denotes that if the anomaly degree of time slot $r$ is more serious than the one of time slot $k$, the optimal collated anomaly score of time slot $r$ is bigger than that of time slot $k$. Property 2 denotes that if the difference in anomaly severity between time slots $r$ and $k$ is greater than the difference between time slots $r^{\prime}$ and $k^{\prime}$, the difference in optimal collated anomaly scores will accurately reflect this gap.
@@
3.4 Train and Inference
During the training process, we use $\mathcal{L}_{a}+\mathcal{L}_{\beta}$ as the loss function. During the inference process, we use the collated anomaly score output by the conditional network as the anomaly score. The Peaks-over-Threshold (POT) algorithm (Siffer et al., 2017) can be used to determine the decision threshold. When the anomaly score is higher than the threshold, the time slot is judged as an anomaly.
@@ 4.1 Experiment Setup
Baselines. We compare CoLLaTe with state-of-the-art (SOTA) anomaly detection methods, including DCdetector (Yang et al., 2023), AnomalyTransformer (Xu et al., 2022), TranAD (Tuli et al., 2022), OmniAnomaly (Su et al., 2019), and MSCRED (Zhang et al., 2019). Additionally, we evaluate against SOTA LLM-based anomaly detection methods, including GPT-4, LLMAD (Liu et al., 2024b), and sigLLM (Alnegheimish et al., 2024). To better understand the contributions of each component in CoLLaTe, we conduct ablation studies by introducing several variations of the model. Specifically, CoLLaTe refers to the version of CoLLaTe without the alignment module. CoLLaTe replaces the collaborative loss with MSE. CoLLaTe modifies the weights $\frac{D_{inter}}{D_{inter}+D_{intra}}$ and $\frac{D_{intra}}{D_{inter}+D_{intra}}$ to $1$. Lastly, CoLLaTe removes the LLM and the conditional network entirely.
Datasets. To verify CoLLaTe can effectively embed the expertise knowledge, we choose three datasets for a complex system, which has domain-specific rules for anomaly detection and highly rely on expert knowledge. One is cluster-wide slowdown detection in cloud environment (Amvrosiadis et al., 2018; Chen et al., 2024b) and two are aircraft monitoring datasets collected from an aircraft manufacturing company that is one of the world’s top 150 enterprises. Additionally, we included a subtle mathematical-assumption-based anomaly detection dataset MGAB (Thill et al., 2020), in which anomalies can not be recognized visually, but need complex mathematical derivation and verification. For more details, please refer to Appendix A.6.
@@
Evaluation Metrics. We use three of the most popular evaluation metrics to measure the performance of CoLLaTe, as many works did (Chen et al., 2022b, 2021; Li et al., 2023): the recall, precision and F1 score.

### v4

@@ 1 Introduction
Recently, methods based on Large Language Models (LLMs) have demonstrated the capablility of reading professional documents, effectively acquiring human knowledge to guide relevant tasks. However, they are often insensitive to the value fluctuations in time series data, and their natural language processing-based representations do not align well with the characteristics of time series data Jin et al., (2024). In contrast, task-specific small scaled models, such as task-specific anomaly detection models (TSADM), typically lack the capabilities of reading professional documents and extracting expert knowledge. However, these models are specifically designed to learn patterns for anomaly detection and often exhibit superior performance when applied to well-matched anomaly detection datasets Zhou et al., (2023). Despite their strengths, TSADMs also have notable limitations. First, for complex systems, such as cloud service system and flight systems, which require lots of expert knowledge to detect and diagnose anomalies accurately, the anomaly assumptions of unsupervised anomaly detection methods can be inconsistent with these scenarios Zha et al., (2020). For instance, anomalies identified based on the voting mechanism of redundant aircraft channels Rice and Mccorkle, (1979) may differ from the outlier assumptions utilized by unsupervised anomaly detection methods Blázquez-García et al., (2021). Thus, researchers need to modify TSADMs to incorporate the background expertise to achieve optimal performance. For instance, anomaly detection methods tailored for cloud service monitoring Ma et al., (2021); Chen et al., 2024b or aircraft monitoring e Silva and Murcca, (2023) have been modified to suit these specific contexts. Second, in many practical scenarios, such as aircraft monitoring Nanduri and Sherry, (2016), it is not feasible to collect system monitoring data for all possible flight conditions. These conditions often represent distinct distributions and normal patterns in the monitoring data. This insufficient coverage of comprehensive data poses a significant challenge, greatly impeding the performance of TSADM.
@@
(a) The performance of CoLLaTe, LLM and TSADM
(b) Misalignment between anomaly scores of the LLM and TSADM
(c) Anomaly score distribution of large language model
(d) Misalignment between original anomaly scores of LLM and TSADM
Figure 1: (a) The anomaly scores of LLM, TSADM, and CoLLaTe. (b) The normalized anomaly score distribution of TSADM and the fitted curve of anomaly scores from the LLM on Mustang Amvrosiadis et al., (2018). (c) shows the histogram of the anomaly score of LLM on the Mustang Amvrosiadis et al., (2018) dataset and the fitted curve of the anomaly score. (d) The original anomaly score distribution of TSADM for anomaly detection and the fitted curve of anomaly scores from the LLM on Mustang.
@@ 2.1 Overview
The architecture of CoLLaTe is depicted in Fig. 2. The anomaly scores, $\dot{s}$ and $S$, are generated by a TSADM and an LLM, respectively, where a set-up pitch prompt is utilized to embed expert knowledge from professional document to the LLM (Appendix. A.12). The TSADM is an anomaly attention Xu et al., (2022) based anomaly detection model, whose architecture is illustrated in Appendix. A.10 in detail. These anomaly scores, $\dot{s}$ and $S$, typically represent different score interpretations, as the LLM and the TSADM can assign the same score to denote varying degrees of anomaly. To address this, we introduce an alignment module to align the different interpretations of the anomaly scores between the LLM and the TSADM. Following alignment, we employ a conditional network to synthesize the aligned scores from the TSADM and anomaly scores from the LLM, using the data representation $R$ as a condition. This process yields the collated anomaly score, $\hat{S}$, which is deemed as the final anomaly score to distinguish anomaly. Additionally, we propose a collaborative loss function that effectively leverages the complementary strengths of the LLM and the TSADM. Notably, we prove this loss function can mitigate error accumulation during the collaborative process between the LLM and the TSADM mathematically and experimentally.
@@ 2.2 Alignment Module
TSADM and large language models (LLMs) often interpret anomaly scores differently. In other words, they may assign the same anomaly score to represent varying degrees of anomaly severity. This discrepancy arises due to two main factors: value range misalignment and distribution misalignment. The value range misalignment is shown in Fig. 1(d), where we shift the fitted curve of LLM anomaly scores to the histogram of TSADM anomaly scores. Most reconstruction-based TSADMs use reconstruction error as the anomaly score, resulting in a highly variable range of scores. In contrast, the anomaly scores produced by LLMs are typically constrained to the range $[0,1]$. This problem can be easily solved by min-max normalization and we obtain the scaled anomaly scores of TSADM, $s$. This process is illustrated in detail in Appendix. A.13.
Even after scaling the anomaly scores of the TSADM, differences in the distribution of anomaly scores can still lead to inconsistent interpretations. As shown in Fig. 1(c), we use a half-Gaussian distribution, which will be elaborated in the following, to fit the histogram of LLM anomaly scores. Then, we shift the fitted curve to the histogram of the anomaly scores from the TSADM in Fig. 1(b), where we observe that most LLM anomaly scores fall below 0.3, whereas the TSADM’s anomaly scores are concentrated in the range of [0.3, 0.4]. As a result, an anomaly score of 0.4 might indicate moderate anomaly risks for the LLM, while the TSADM may interpret it as normal. In such cases, direct collaboration between the TSADM and the LLM without proper alignment could misrepresent their intended outputs. For example, if the LLM assigns a score of 0.4 to a time slot, the TSADM might incorrectly assume the LLM considers the time slot normal. To address this challenge, we propose an Alignment Module to align their interpretations and ensure effective collaboration.
@@
Subsequently, we use a mapping function $\mathcal{M}(s)$ to map $s\in\{s_{i}\}_{i=0}^{n}$ from the TSADM to a new set, whose distribution should approach $f(S)$. To achieve this, we divide the value range of anomaly scores, $[0,1]$, into N bins. We use $c_{i}$ to denote how many mapped scores $\mathcal{M}(s)$ fall into the $i^{th}$ bin. Then, we minimize cross entropy to decrease the distance between $f(S)$ and the distribution mapped set as Eq. 2 shows, where $\hat{\mu},\hat{\sigma}$ are the mean and standard variance of distribution $f(S)$, $\mu_{M}=\frac{1}{n}\sum_{k=1}^{n}\mathcal{M}(s_{k})$, and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are hyperparameters and bigger than zero. One problem with Eq. 2 is that the counting process of $c_{i}$ is non-differentiable and could not be used in backforward propagation.
@@
To solve the above problem, we prove that when $N$ approaches infinite, the objective function in Eq. 2 is equivalent to $-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}$, which is differentiable. The proof is given in Appendix. A.1. Consequently, we use Eq. 3 as our loss function.
$\min_{\mathcal{M}}.\mathcal{L}_{a}=-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}\\ +\hat{\lambda}_{2}(\frac{1}{n-1}\sum_{i=1}^{n}(\mathcal{M}(s_{i})-\mu_{M})^{2}-\hat{\sigma}^{2})^{2}$ (3)
@@ 2.3 Collaborative Loss
Many anomaly detection studies categorize anomalies into contextual anomalies and point anomalies Schmidl et al., (2022); Tuli et al., (2022). We conducted experiments on two aircraft monitoring datasets, which are detailed in Appendix A.7. The results, presented in Tab. 1, demonstrate the complementary performance of the TSADM and the LLM in detecting contextual and point anomalies. Specifically, the F1 score for contextual anomalies is higher for the TSADM compared to GPT-4, while GPT-4 achieves a better F1 score for point anomalies than the TSADM. According to this insight, we propose a collaborative loss function that adaptively optimizes the collated anomaly score. Specifically, we begin by dividing the time series into patches Nie et al., (2023). For each time slot, we compute the average distance between the present time slot and all other time slots within the same patch, denoted as $D_{intra}$. Next, we calculate the average distance between the patch of the current time slot and all other patches, denoted as $D_{inter}$. Since anomalous time slots differ significantly from normal ones, while normal time slots tend to share similarities, the metrics $D_{intra}$ and $D_{inter}$ behave differently for point and contextual anomalies. For point anomalies, $D_{intra}$ should be large because the point anomaly stands out distinctly from the surrounding normal time slots. Conversely, $D_{inter}$ for point anomalies should be smaller, as the normal time slots within the current patch resemble those in other patches, reducing the average difference. For contextual anomalies, which typically involve a sustained deviation, the anomalies within an anomalous event are similar to each other but differ significantly from normal time slots. As a result, $D_{intra}$ should be small, while $D_{inter}$ should be large. Considering LLM can handle point anomaly better and TSADM can handle contextual anomaly better, we design the collaborative loss function in Eq. 5, where $\mathcal{L}(a,b)$ represents a loss function that measures the distance between $a$ and $b$.
@@
Assumption 3. $\mathcal{L}_{\beta}^{*}$ in Eq. 6 is $L$-Lipschitz continuous.
@@
We examined the validity of these Assumptions in the context of our experiment through empirical discussion, which is shown in Appendix. A.17.
@@
From Theorem 1, we observe that when MSE is used as $\mathcal{L}$, the expected prediction error of the conditional network exceeds the weighted sum of the expected prediction errors of the TSADM and LLM. Given that the weights $\lambda_{1}$ and $\lambda_{2}$ are both greater than 0 and sum to 1, it indicates that the collaborative error, when using MSE as $L_{(a,b)}$, consistently exceeds the smaller of the errors produced by the LLM and TSADM. This phenomenon reflects error accumulation and is further validated experimentally, as discussed in detail in Section 3.4.
To solve this problem we propose the collaborative loss function as shown in Eq. 7, where $\hat{S}_{i}$, $s_{i}$ and $S_{i}$ are anomaly scores of $i^{t}h$ sample output by the conditional network, TSADM and LLM respectively. In Lemma 1, we prove that using the collaborative loss function is approximate to using the loss function in Eq. 6. This equivalence shows that Eq. 7 prevents error accumulation between the TSADM and LLM, as Eq.6 directly minimizes the difference between the collated anomaly scores of two samples and the ground truth. Additionally, we experimentally confirm that Eq. 7 effectively avoids error accumulation, as discussed in Section. 3.4.
@@
Moreover, in Theorem 2, we prove that the optimal solution of Eq. 6 satisfies two key properties. Property 1 denotes that if the anomaly degree of time slot $r$ is more serious than the one of time slot $k$, the optimal collated anomaly score of time slot $r$ is bigger than that of time slot $k$. Property 2 denotes that if the difference in anomaly severity between time slots $r$ and $k$ is greater than the difference between time slots $r^{\prime}$ and $k^{\prime}$, the difference in optimal collated anomaly scores will accurately reflect this gap.
@@ 3.1 Experiment Setup
Baselines. We compare CoLLaTe with state-of-the-art (SOTA) anomaly detection methods, including DCdetector Yang et al., (2023), AnomalyTransformer Xu et al., (2022), TranAD Tuli et al., (2022), OmniAnomaly Su et al., (2019), and MSCRED Zhang et al., (2019). Additionally, we evaluate against SOTA LLM-based anomaly detection methods, including GPT-4, LLMAD Liu et al., 2024b , and sigLLM Alnegheimish et al., (2024). To better understand the contributions of each component in CoLLaTe, we conduct ablation studies by introducing several variations of the model. Specifically, CoLLaTe refers to the version of CoLLaTe without the alignment module. CoLLaTe replaces the collaborative loss with MSE. CoLLaTe modifies the weights $\lambda_{1}$ and $\lambda_{2}$ to $1$. Lastly, CoLLaTe removes the LLM and the conditional network entirely.
Datasets. To verify CoLLaTe can effectively embed the expertise knowledge, we choose three datasets for a complex system, which has domain-specific rules for anomaly detection and highly rely on expert knowledge. One is cluster-wide slowdown detection in cloud environment Amvrosiadis et al., (2018); Chen et al., 2024b and two are aircraft monitoring datasets collected from an aircraft manufacturing company that is one of the world’s top 150 enterprises. Additionally, we included a subtle mathematical-assumption-based anomaly detection dataset MGAB Thill et al., (2020), in which anomalies can not be recognized visually, but need complex mathematical derivation and verification. For more details, please refer to Appendix A.7.
@@
Evaluation Metrics. We use three of the most popular evaluation metrics to measure the performance of CoLLaTe, as many works did Chen et al., 2022b ; Chen et al., (2021); Li et al., (2023): the recall, precision and F1 score.